Metallocene are metals sandwiched by two cyclopentadienyl ions, (C5H5)–. A metallocene sample of unknown metal, MC5H5 (all 5 are subscripts) was burned in air. Products of combustion found to contain 46.08488% carbon dioxide, 9.42644% water vapour and the rest are metal (I) oxide. All products are in mass percent. The mass ratio of stoichiometric amount of oxygen to the burned metallocene is 1:1.295. What is the name of the metal in the metallocene sample? Prove by showing your complete solution.
1
Expert's answer
2019-08-14T02:43:41-0400

The reaction of combustion of the given metallocene (MeCp) is following:

"2 \ MeC_5H_5 + 13O_2=10CO_2+5H_2O+Me_2O"

The stoichiometric ratio between MeCp and O2 is:

"\frac{n(MeCp)}{n(O_2)}=\frac{2}{13}"

As "n(MeCp)=\frac{m(MeCp)}{M(MeCp)}" and "n(O_2)=\frac{m(O_2)}{M(O_2)}" ,the above equation transforms to the following:

"\frac{m(MeCp)}{m(O_2)}\cdot \frac{M(O_2)}{M(MeCp)}=\frac{2}{13}"

From the given conditions the mass ratio of stoichiometric amount of oxygen to the burned metallocene is 1:1.295, i.e. "\frac{m(O_2)}{m(MeCp)}=\frac{1}{1.295}"

Thus two equations above transform to the following:

"\frac{1.295}{1}\cdot \frac{M(O_2)}{M(MeCp)}=\frac{2}{13}"

from which we can find molar mass of metallocene:

"M(MeCp)=\frac{1.295\cdot 13 \cdot 32\frac{g}{mol}}{2}=269.36\frac{g}{mol}"

Taking into account that "M(MeCp)=M(Me)+M(Cp)=M(Me)+M(C_5H_5)=M(Me)+65 \ (\frac{g}{mol})" it is easy to find the molar mass of metal:

"M(Me)+65=269.36"

"M(Me)=204.36 \frac{g}{mol}"

This molar mass of metal corresponds to thallium, that is known to form a metallocene that adopts a sandwich-like polymeric structure: